OpenStudy (anonymous):

Find the value of the the constant k for which the equation gf(x)= k has equal roots. f:x |--> 4x-2x^2 g:x |--> 5x+3

OpenStudy (anonymous):

is it g(f(x)) = k?

OpenStudy (anonymous):

No. It's gf(x). Which is 5(4x-2x^2)+3

OpenStudy (kropot72):

Using the standard formula for finding the roots of a quadratic and substituting with C being the required constant for equal roots the following term must equal zero. \[\sqrt{(400}+40C)=0\] Therefore the required constant C=-10

OpenStudy (anonymous):

I thought you used b-4ac? And that's not the answer...?

OpenStudy (anonymous):

But how do you plug in the values into b-4ac?

OpenStudy (kropot72):

The roots are 1,1

OpenStudy (anonymous):

i got k=13. do you know the answer? i think if finally understand the question..

OpenStudy (kropot72):

I agree with dpalnc. I forgot to do the last step:(

OpenStudy (kropot72):

I used: \[b ^{2}-4ac\]

OpenStudy (anonymous):

How? I don't know how to plug in the values..?

OpenStudy (kropot72):

I first rearranged as follows: \[-10x ^{2}+20x+C\] Then a=-10, b=20 and c=C In that case: \[b ^{2}-4ac=400-4(-10*C)=0\] when C=-10 Therefore the required constant on the RHS of the original equation must be 13. Subtracting 13 from both sides makes the quadratic equal to zero and the constant term becomes -10.

OpenStudy (anonymous):

That doesn't seem to make much sense. C is already given... =3, and the workings don't match..

OpenStudy (kropot72):

I ignored the given constant 3 and concentrated on finding what constant is required to form a quadratic equation using the two given variable terms. Having found what constant was needed it was simple to find the value of k (13) which gives a quadratic with equal roots when shifted to the LHS. 3-13=10.

OpenStudy (kropot72):

\[-10x ^{2}+20x-10=(-5x+5)(2x-2)\] The roots are 1,1

OpenStudy (kropot72):

Sorry for the typo above. 3-13=-10

OpenStudy (kropot72):

WolframAlpha indicates that k=13 will give roots of 1,1. Can anyone see something wrong with my method of arriving at the same result?

OpenStudy (anonymous):

is 13 the right answer? 13 is what i got with my interpretation of the problem

OpenStudy (anonymous):

 equal roots mean the discriminant = 0 (double root) so just set b^2-4ac = 0.